OpenStudy (anonymous):

how to graph x+y<3 y>x+1

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

rearrange x+y<3 to get y<-x+3 now draw graph for y=-x+3 and graph for y=x+1 now shade the region under y<-x+3 and shade region over y>x+1. that will be the inequalities drawn and relevant regions shaded

OpenStudy (anonymous):

thank you